OpenStudy (some.random.cool.kid):

Which system of equations is represented by the graph? a line is graphed through points negative 6 comma 7 and negative 1 comma 2. It intersects a parabola that opens up at these two points. y = x2 − 6x − 7 x − y = 1 y = x2 − 6x + 7 x + y = −1 y = x2 + 6x − 7 x − y = 1 y = x2 + 6x + 7 x + y = 1 @pooja195 @jabez177 @zepdriz @mathmale @whpalmer4 @sammixboo @.sam. @AloneS

OpenStudy (some.random.cool.kid):

@s

OpenStudy (some.random.cool.kid):

@zepdrix

OpenStudy (some.random.cool.kid):

@sweetburger

OpenStudy (some.random.cool.kid):

y = x2 − 6x − 7 x − y = 1 y = x2 − 6x + 7 x + y = −1 y = x2 + 6x − 7 x − y = 1 y = x2 + 6x + 7 x + y = 1

OpenStudy (some.random.cool.kid):

@zzr0ck3r @zepdrix ... you guys are math elites right?

OpenStudy (zzr0ck3r):

can you figure out the equation of the straight line?

OpenStudy (zzr0ck3r):

@some.random.cool.kid

OpenStudy (some.random.cool.kid):

i am here

OpenStudy (some.random.cool.kid):

I have little idea how to do this though... Mind showing me an example of how i tdone then I try one based off of your template?

OpenStudy (some.random.cool.kid):

@zzr0ck3r

OpenStudy (zzr0ck3r):

Do you know what the slope of a line is?

OpenStudy (zzr0ck3r):

\(\dfrac{\text{RISE}}{\text{RUN}}\)

OpenStudy (some.random.cool.kid):

yes

OpenStudy (some.random.cool.kid):

ill brb

OpenStudy (radar):